OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 8pi

OpenStudy (bahrom7893):

Terri u asked at least 15 of these.. C = 2*pi*r A = pi*r^2

OpenStudy (bahrom7893):

2*pi*r = 8pi so r = 4 now plug that into the A formula

OpenStudy (bahrom7893):

what do u get?

OpenStudy (anonymous):

25.15

OpenStudy (bahrom7893):

Without a calculator, it says find the exact area, write out how you would substitute, please.

OpenStudy (bahrom7893):

A = pi*r^2 r = 4 come on, think..

OpenStudy (bahrom7893):

TERRI! Hello??

OpenStudy (anonymous):

Not 15, more like 50...

OpenStudy (anonymous):

stop treating me like im stupid i need hel its nt my fault that i suck at geometry but if u dnt wanna hep and keepmocking me abt how many questions ipost thenim delting this account

OpenStudy (bahrom7893):

What did I say wrong? I said i've helped u with plenty of these, time u tried one on your own. I can make an excuse that i'm stupid in like chemistry and not do my work..

OpenStudy (anonymous):

That's up to u, it would be nice to see u make some progress... I saw u posted a slightly more interesting question which I answered fully without making any fun of you. But unless u have these basics down u will never be able to answer such questions by yourself.